ECH 4323: Chemical Process Control EXPERIMENTAL PROTOTYPE

Design and develop a table-top windmill that can pump water from one reservoir to another. The project is completely open-ended and any working prototype will be accepted. A starting point for this project may be to do a search on the Archimedes screw pump driven by wind power.

The prototype must be built from commonly available parts. You are not allowed to buy a kit from a hobby shop.

This is a team assignment and will be graded as follows:

Working model  3 points
Engineering and Cost Analysis 2 points
A feature that ”wows” me  3 bonus points